# Order Matters in Hallucination: Reasoning Order as Benchmark and Reflexive Prompting for Large-Language-Models

### v2

@@ Introduction
Although recent studies have demonstrated the impact of prompting strategies on mitigating hallucination problems across various fields (Si et al. 2022; Mündler et al. 2023; Ji et al. 2023; Dhuliawala et al. 2023), there is a lack of research and discussion on the gap between answer-first prompt and logic-first prompt. To fill this gap, we designed a new benchmark method for testing language models based on this observation: Reasoning Order as Benchmark. This method reflects the consistency of the model’s reasoning logic by comparing the differences between the answer-first prompt and the logic-first prompt. Inspired from Chain-of-Thought prompting (Wei et al. 2022) which enhances LLM reasoning performance by adding step-by-step guidance, we additionally designed Reflexive Prompting, a two-step prompting strategy. In the first step, we use both the answer-first prompt and the logic-first prompt to obtain two potentially different answers. In the second step, we analyze these two answers to derive the final answer. Experimental results demonstrate that this method improves the accuracy of reasoning tasks across different datasets and various LLMs. Furthermore, the accuracies achieved using different prompting strategies show strong correlation with the results of the Reasoning Order as Benchmark test result, proving the usability of this benchmarking approach.
To summarize, our contributions are threefold:
@@
  • We provide empirical evidence demonstrating the effectiveness of our methods across multiple datasets and LLM architectures.
@@
Chain-of-Verification (Dhuliawala et al. 2023) developed a Chain-of-Thought alike sequential approach namely Chain-of-Verification: for a given initial draft response, first plans verification questions to check its work, and then systematically answer those questions as a Chain-of-Thought procedure to gradually guide the model and finally produce an improved revised response.
@@ Methods
We noticed that in every response from the large language models with incorrect answer, the first sentence always presents the answer. Considering the sequential nature of language model outputs, at the point of generating the answer, the model has not yet seen the subsequent reasoning portion. Therefore, it may provide an incorrect answer based on similar causes identified in the training data, and then continue to generate reasoning to support this answer. To explore this phenomenon, we designed a set of new prompts for the question “Is 9.11 $>$ 9.9?”: (1) Prompts that instruct the model to first output the final result and then provide the reasoning, and (2) prompts that instruct the model to first output the reasoning and then the corresponding result. Surprisingly, we found that even if the LLMs can give out correct answer under original prompt, the issue of incorrect judgement may still occur when asked to generate answer before reasoning logic, as the comparison using Claude shown in Figure 2.
@@
According to the aforementioned hypothesis, if language models could explicitly access information about potential future outputs during the generation process, modifying the architecture of the decoder could potentially alleviate this issue. Unfortunately, due to a lack of sufficient computational resources and corpora, we are unable to retrain such LLMs. Consequently, we have approached the problem from the perspective of prompt engineering, designing a novel prompting method that enables large language models to acquire this information: Reflexive Prompting.
In reflexive prompting, the query process of LLMs is transformed from a single-step direct inquiry into a two-step procedure. Similarly, in the first step, we generate the “answer first” and “logic first” prompt variants for original query and get the corresponding results. In the second step, the original question and the two results are fed into the LLM altogether as reflexive prompt for the final decision. The query process can be illustrated in Algorithm 2. One example of the reflexive prompt is shown below:
@@
The feasibility of this strategy primarily stems from two aspects. Firstly, providing language models with explicit subsequent logical reasoning information, as posited by prior hypotheses, allows the models to reassess their reasoning processes from multiple perspectives. This process is somewhat analogous to chain-of-thought prompting, which also utilizes the intermediate information output by the model itself to adjust the reasoning process. However, unlike chain-of-thought prompting, which requires human experts to decompose problems into sub-problems to facilitate step-by-step resolution by the model, the method described in this paper does not necessitate such decomposition, thereby enhancing the potential for generalization. Additionally, the variation in results due to the different sequences of reasoning and answers can be viewed, in a sense, as an ensemble learning approach. This involves the model generating potentially divergent responses, which are then stacked by the model itself acting as a meta-model. The inherent reasoning capabilities of LLMs are crucial in this procedure, enabling the employment of self consistency review as meta-model decision makers.
@@
To evaluate the effectiveness of the reasoning order as benchmark as well as the reflexive prompting strategy, we designed the following experiments testing a various of reasoning datasets on different LLMs. Experiment codes and results are available in https://github.com/XieZikai/ReflexivePrompting.
Claude Raw Prompt Answer First Logic First Reflexive Prompt
LogiQA 0.730 0.725 0.724 0.734
TruthfulQA 0.831 0.834 0.826 0.845
MMLU 0.862 0.838 0.860 0.855
Gemini Raw Prompt Answer First Logic First Reflexive Prompt
LogiQA 0.616 0.604 0.622 0.625
TruthfulQA 0.729 0.718 0.735 0.757
MMLU 0.717 0.714 0.769 0.769
GPT Raw Prompt Answer First Logic First Reflexive Prompt
LogiQA 0.624 0.627 0.612 0.627
TruthfulQA 0.623 0.636 0.613 0.639
MMLU 0.811 0.748 0.810 0.815
Llama Raw Prompt Answer First Logic First Reflexive Prompt
LogiQA 0.661 0.680 0.655 0.684
TruthfulQA 0.659 0.690 0.656 0.726
MMLU 0.814 0.764 0.821 0.843
@@
LogiQA TruthfulQA MMLU
Claude 0.870 0.897 0.884
Gemini 0.658 0.807 0.786
Llama 0.759 0.744 0.814
GPT 0.781 0.791 0.826
@@
Pearson Raw Prompt Answer First Logic First Reflexive Prompt
LogiQA 0.906 0.941 0.835 0.884
TruthfulQA 0.977 0.999 0.969 0.981
MMLU 0.903 0.999 0.974 0.873
@@
Datasets The core of the two methods introduced in this paper lies in comparing the outputs generated from the “order” prompt pair, making reasoning tasks particularly suitable for this experiment. Furthermore, multiple-choice datasets provide distinct options as criteria for determining differences, hence, we selected three different multiple-choice reasoning datasets as test sets.
@@
  • Measureing Masssive Multitask Language Understanding(MMLU) (Hendrycks et al. 2020) is a multitask reasoning dataset, covering 57 tasks including elementary mathematics, US history, computer science, law and more. The validation set is chosen for experiments, which contains 1,531 multiple-choice questions.
@@
  • TruthfulQA (Lin, Hilton, and Evans 2021) is a benchmark to measure whether a language model is truthful in generating answers that humans would answer falsely due to a false belief or misconception. The dataset comprises 817 questions that span 38 categories, including health, law, finance and politics.
@@
  • LogiQA (Liu et al. 2020) is an expert-written question dataset for testing human logical reasoning, which consists of 8,678 QA instances, covering multiple types of deductive reasoning. Due to limited budget, only the first 1,000 questions are used in the experiments.
@@
Models To conduct our experiments, we selected four commonly used large language models available on the market as test subjects, employing unadorned raw prompts, answer-first prompts, and logic-first prompts as baselines. These models are: GPT-4o-mini (Achiam et al. 2023), Llama-3.1-70b (Vavekanand and Sam 2024), Claude-3.5-sonnet (Amazon 2024) and Gemini-1.5-flash (Reid et al. 2024). It is important to note that since chain-of-thought prompting requires the design of specialized decomposition processes for each question, effectively introducing additional information, we consider it inappropriate to compare it directly with the other approaches.
@@
Experimental Results
We use the accuracy for multiple-choice questions as comparison criteria among raw prompts, answer-first prompts, logic-first prompts and reflexive prompts. The results for all the LLMs and datasets are shown below in Table 1.
Given that this method is based on the phenomenon of inconsistent outputs from the answer-first prompt and logic-first prompt, it is natural to examine the extent of these inconsistencies in the model outputs, which constitutes the key methodology of Reasoning Order as Benchmark. The consistency of all models on each dataset is shown in Table 3. To further check the statistical properties, we calculated the Pearson correlation coefficient between the consistency and the accuracy of all models on each dataset and each prompting strategy, which is shown in Table 3. We can observe a very clear linear correlation between the models’ consistency and accuracy. This is likely because low consistency may result from the models’ inability to output answers with high confidence, indicating that the model has not learned enough to produce reliable answers, thereby leading to low accuracy. This result indicates that Reasoning Order as Benchmark can indeed serve as a meaningful benchmark indicator.
@@
In this paper, we delve into the potential causes of the “9.11$>$9.9” problem and introduce the methods of Reasoning Order as Benchmark and Reflexive Prompting. By assigning both answer-first and logic-first prompts to the same question, we observed inconsistencies in the models’ responses depending on the order of answer generation and logical reasoning. To drive this issue, we designed the reflexive prompt to return the inconsistent answer-logic output pairs back to the model simultaneously as a step-two evaluation and these inconsistencies are mitigated to some extent, thereby enhancing the reasoning performance. Evaluation on the benchmark method shows high linear relationship with model accuracy, implicating that additional module designed for reasoning order consistency may help boosting model performance.
It must be acknowledged that this work still possesses several limitations. Firstly, due to budget constraints, we were unable to conduct experiments on larger datasets or with more extensive large language models. Although we believe that the current experiments sufficiently demonstrate the viability and performance boost of the prompt strategy introduced in this research, additional experiments could potentially reveal more reliable patterns from the benchmark results. Furthermore, we lacked the computational resources and data necessary to retrain large language models. Originating from the hypotheses proposed in Section LABEL:error_reproduction, we posit that improvements could be made to large language models by enhancing the decoder part. Since the current output cannot see subsequent outputs, a potential research direction involves integrating the output layer with the original hidden state vectors before output layer, enabling the decoder to “see” the previous response during output. (Goyal et al. 2023) has already shown the potential to pause token to apply inference-time delay, introducing performance gain on various of reasoning datasets, showing that allowing for revisiting and revising its own answers may mitigate such issues. Therefore, a modified decoder network to explicitly combine possible future output may be a direct approach to mitigate hallucination in language models.

### v3

@@ Introduction
Although recent studies have demonstrated the impact of prompting strategies on mitigating hallucination problems across various fields (Si et al. 2022; Mündler et al. 2023; Ji et al. 2023; Dhuliawala et al. 2023), there is a lack of research and discussion on the gap between answer-first prompt and logic-first prompt. To fill this gap, we designed a new benchmark method for testing language models based on this observation: Reasoning Order as Benchmark. This method reflects the consistency of the model’s reasoning logic by comparing the differences between the answer-first prompt and the logic-first prompt. Inspired from Chain-of-Thought (CoT) prompting (Wei et al. 2022) which enhances the performance of LLM reasoning by adding step-by-step guidance, we also designed Reflexive Prompting, a two-step prompting strategy. In the first step, we use both the answer-first prompt and the logic-first prompt to obtain two potentially different answers. In the second step, we analyze these two answers to derive the final answer. Experimental results demonstrate that this method improves the accuracy of reasoning tasks across different datasets and various LLMs. Furthermore, the accuracies achieved using different prompting strategies show strong correlation with the results of the Reasoning Order as Benchmark test result, proving the usability of this benchmarking approach.
In summary, our contributions are threefold:
@@
  • We provide empirical evidence that demonstrates the effectiveness of our methods in multiple datasets and LLM architectures.
@@
Chain-of-Verification (Dhuliawala et al. 2023) developed a CoT alike sequential approach namely Chain-of-Verification: for a given initial draft response, first plans verification questions to check its work, and then systematically answer those questions as a CoT procedure to gradually guide the model and finally produce an improved revised response.
@@ Methods
We noticed that in every response from the large language models with incorrect answer, the first sentence always presents the answer. Considering the sequential nature of language model outputs, at the point of generating the answer, the model has not yet seen the subsequent reasoning portion. Therefore, it may provide an incorrect answer based on similar causes identified in the training data, and then continue to generate reasoning to support this answer. To explore this phenomenon, we designed a set of new prompts for the question “Is 9.11 $>$ 9.9?”: (1) Prompts that instruct the model to first output the final result and then provide the reasoning, and (2) prompts that instruct the model to first output the reasoning and then the corresponding result. Surprisingly, we found that even if the LLMs can give out correct answer under original prompt, the issue of incorrect judgement may still occur when asked to generate an answer before reasoning logic, as in the comparison using Claude shown in Figure 2. Based on such observation, we propose the following hypotheses:
  • The sequential output of LLMs may lack of considerations for following outputs, sometimes cause hallucinations based on wrong pre-posing context;
  • Answer-first and logic-first prompt pair may efficiently trigger such hallucinations;
  • By manually providing such hallucination information, LLMs can revise the inference process and choose the correct answer.
@@
Clearly, a robust language model should behave like a human, producing the same output for paraphrased queries that do not alter the underlying meaning. Consequently, evaluating the semantic stability of the model under such variations can indirectly validate its accuracy and robustness. This testing procedure is somewhat analogous to zero-shot adversarial attack-based evaluation methods (Shayegani et al. 2023), but it holds an advantage derived from Hypothesis 2: a reflexive prompt pair triggers adversarial attacks more efficiently. By sacrificing a portion of accuracy, this approach reduces testing time and cost, as it does not require additional computational resources to generate adversarial examples.
@@
According to the aforementioned hypothesis, if language models could explicitly access information about potential future outputs during the generation process, modifying the architecture of the decoder could potentially alleviate this issue. Unfortunately, due to the lack of sufficient computational resources and corpora, we are unable to retrain such LLMs. Consequently, we have approached the problem from the perspective of prompt engineering, designing a novel prompting method that enables large language models to acquire this information: Reflexive Prompting.
In reflexive prompting, the LLM query process is transformed from a single-step direct inquiry into a two-step procedure. Similarly, in the first step, we generate the “answer first” and “logic first” prompt variants for the original query and get the corresponding results. In the second step, the original question and the two results are fed into the LLM altogether as a reflexive prompt for the final decision. The query process can be illustrated in Algorithm 2. An example of a reflexive prompt is shown below:
@@
The feasibility of this strategy primarily stems from two aspects. First, providing language models with explicit subsequent logical reasoning information, as posited by prior hypotheses, allows the models to reassess their reasoning processes from multiple perspectives. This process is somewhat analogous to CoT prompting, which also utilizes the intermediate information output by the model itself to adjust the reasoning process. However, unlike CoT prompting, which requires human experts to decompose problems into subproblems to facilitate step-by-step resolution by the model, the method described in this paper does not necessitate such decomposition, thereby enhancing the potential for generalization. Relatively, the logic-first prompt can be considered as a zero-shot CoT prompt since the answer is generated “step by step”, from which we derive our more efficient prompt strategy. Additionally, the variation in results due to the different sequences of reasoning and answers can be viewed, in a sense, as an ensemble learning approach. This involves the model generating potentially divergent responses, which are then stacked by the model itself acting as a meta-model. The inherent reasoning capabilities of LLMs are crucial in this procedure, enabling the employment of self-consistency review as meta-model decision makers.
@@
To evaluate the effectiveness of the reasoning order as benchmark as well as the reflexive prompting strategy, we designed the following experiments testing a variety of reasoning datasets on different LLMs. Experiment codes and results are available in https://github.com/XieZikai/ReflexivePrompting.
Claude Raw Prompt Answer First Logic First (0-shot CoT) Answer Ensemble Reflexive Prompt
LogiQA 0.730 0.725 0.724 0.733 0.734
TruthfulQA 0.831 0.834 0.826 0.843 0.845
MMLU 0.862 0.838 0.860 0.866 0.855
BigBench 0.871 0.888 0.888 0.890 0.895
Gemini Raw Prompt Answer First Logic First (0-shot CoT) Answer Ensemble Reflexive Prompt
LogiQA 0.616 0.604 0.622 0.643 0.625
TruthfulQA 0.729 0.718 0.734 0.736 0.757
MMLU 0.717 0.714 0.769 0.764 0.769
BigBench 0.826 0.832 0.817 0.835 0.835
Llama Raw Prompt Answer First Logic First (0-shot CoT) Answer Ensemble Reflexive Prompt
LogiQA 0.661 0.680 0.655 0.684 0.684
TruthfulQA 0.659 0.690 0.656 0.757 0.726
MMLU 0.814 0.764 0.821 0.832 0.843
BigBench 0.841 0.823 0.859 0.843 0.862
GPT Raw Prompt Answer First Logic First (0-shot CoT) Answer Ensemble Reflexive Prompt
LogiQA 0.624 0.627 0.612 0.620 0.627
TruthfulQA 0.623 0.636 0.613 0.609 0.639
MMLU 0.811 0.748 0.810 0.807 0.815
BigBench 0.862 0.850 0.817 0.866 0.840
@@
LogiQA TruthfulQA MMLU BigBench
Claude 0.870 0.897 0.884 0.902
Gemini 0.658 0.807 0.786 0.830
Llama 0.759 0.744 0.814 0.888
GPT 0.781 0.791 0.826 0.855
@@
Pearson Raw Prompt Answer First Logic First (0-shot CoT) Answer Ensemble Reflexive Prompt
LogiQA 0.906 0.941 0.835 0.664 0.884
TruthfulQA 0.977 0.999 0.969 0.557 0.981
MMLU 0.903 0.999 0.974 0.900 0.873
BigBench 0.650 0.514 0.937 0.644 0.947
@@
Preliminary Experiment
As noted above, one key hypothesis behind the idea of this project is that the answer-first and logic-first prompt pair can efficiently generate contradictory completions, so that the LLM may collect knowledge from both sides and infer the correct one. To verify this hypothesis as a sanity check, we compared our approach with the self-contradiction trigger algorithm proposed in (Mündler et al. 2023) in terms of the efficiency of triggering the contradiction. We follow the same prompt settings used in (Mündler et al. 2023), adding these two extra descriptions to form answer-first and logic-first strategies:
Please give me the logic/answer then give me the answer/logic.
We chose ChatGPT with GPT-3.5 turbo to carry out the experiment. With only 2 trials of prompt generating (answer-first and logic-first strategies), 7 out of 30 test questions triggered contradictions, achieving 23.3% self-contradiction trigger rate, which outperforms 17.7% reported from (Mündler et al. 2023). Although the experiment scale is rather limited and the experimental conditions are not entirely identical, considering that our approach is a noniterative, zero-shot method without additional knowledge, the result still validates the correctness of our hypothesis 2: Answer-first and logic-first prompt pairs may efficiently trigger the hallucination of self-inconsistency.
@@
Datasets The core of the two methods introduced in this paper lies in comparing the outputs generated from the “order” prompt pair, making reasoning tasks particularly suitable for this experiment. Furthermore, multiple choice datasets provide distinct options as criteria for determining differences to trigger potential hallucinations, and hence, we selected four different multiple choice reasoning datasets as test sets. Fill-in-the-blank and short answer question datasets are not included, since the evaluating the differences of the more diverse answers requires extra resource.
@@
  • Measureing Masssive Multitask Language Understanding(MMLU) (Hendrycks et al. 2020) is a multitask reasoning dataset, covering 57 tasks including elementary mathematics, US history, computer science, law and more. The validation set is chosen for the experiments, which contains 1,531 multiple choice questions.
@@
  • TruthfulQA (Lin, Hilton, and Evans 2021) is a benchmark to measure whether a language model is truthful in generating answers that humans would answer falsely due to a false belief or misconception. The dataset comprises 817 questions that span 38 categories, including health, law, finance, and politics.
@@
  • LogiQA (Liu et al. 2020) is an expert-written question dataset for testing human logical reasoning, which consists of 8,678 QA instances, covering multiple types of deductive reasoning. Due to a limited budget, only the first 1,000 questions are used in the experiments.
  • BigBench (bench authors 2023) contains 204 different tasks, covering areas such as linguistics, childhood development, math, common-sense reasoning and more. The multiple choice task “ruin_names” with 447 questions is chosen to evaluate the prompt strategies, which focus on selecting the humorous edit that ’ruins’ the input movie or musical artist name.
@@
Models To conduct our experiments, we selected four commonly used large language models available on the market as test subjects, employing unadorned raw prompts, answer-first prompts, and logic-first prompts as baselines. These models are: GPT-4o-mini (Achiam et al. 2023), Llama-3.1-70b (Vavekanand and Sam 2024), Claude-3.5-sonnet (Amazon 2024) and Gemini-1.5-flash (Reid et al. 2024). In addition, it seems like an unfair comparison between reflexive prompt and the former prompt strategies for the extra LLM evaluations introducing extra model inferences and cost. Therefore, the ensemble model of three strategies voting is also added as a baseline.
Chain-of-thought prompting seems like a natural baseline for comparison. However, the logic-first prompting can be considered as equivalent to zero-shot CoT prompting (please see case study Figure 4 and Figure 4 for detail), and few-shot CoT prompting requires the design of specialized decomposition processes for each question, effectively introducing additional information. Therefore, we consider it unnecessary to compare it directly with few-shot CoT prompting.
Experimental Results
We use the accuracy for multiple choice questions as comparison criteria among raw prompts, answer-first prompts, logic-first prompts, answer ensemble and reflexive prompts. The results for all the LLMs and datasets are shown in Table 3.
Furthermore, we observe that the ensemble model combining three different prompt strategies, while slightly inferior overall to our method, performs only marginally worse. This is because our approach does not introduce any additional knowledge, such as human corrections of erroneous responses or step-by-step answer revisions. Ontologically, both methods exploit the idea of assembling multiple weaker classifiers into a stronger one, and since the number of these weaker classifiers is relatively small, expecting them to vastly surpass each individual classifier on its own is unrealistic. Therefore, even a relatively small performance gap is sufficient to demonstrate the advantages of our method. From this perspective, the strategy in which an LLM re-examines and self-validates its contradictory responses proves to be a more effective ensemble learning scheme.
@@
Given that this method is based on the phenomenon of inconsistent outputs from the answer-first prompt and logic-first prompt, it is natural to examine the extent of these inconsistencies in the model outputs, which constitutes the key methodology of Reasoning Order as Benchmark. The consistency of all models on each dataset is shown in Table 3. To further check the statistical properties, we calculated the Pearson correlation coefficient between the consistency and the accuracy of all models on each dataset and each prompting strategy, which is shown in Table 3. Results with p-value less than 0.05 are shown in bold font. We can observe a very clear linear correlation between the models’ consistency and accuracy in most of the datasets and prompt strategies with low p-values, showing the feasibility of using this benchmark as a cost-efficient preliminary test. This is likely because low consistency may result from the models’ inability to output answers with high confidence, indicating that the model has not learned enough to produce reliable answers, thereby leading to low accuracy. We can also observe a degree of discrepancy between the performance of the answer ensemble and the model’s consistency, stemming from the fact that once ensemble methods are employed, the final answer is no longer a direct output of the model. Nonetheless, the two remain sufficiently correlated, offering a certain level of guidance for model performance. This result indicates that Reasoning Order as Benchmark method can indeed serve as a meaningful benchmark indicator.
@@
In this paper, we delve into the potential causes of the “9.11$>$9.9” problem and introduce the methods of Reasoning Order as Benchmark and Reflexive Prompting strategy. By assigning both answer-first and logic-first prompts to the same question, we observed inconsistencies in the models’ responses depending on the order of answer generation and logical reasoning. To drive this issue, we designed the reflexive prompt to return the inconsistent answer-logic output pairs back to the model simultaneously as a step-two evaluation and these inconsistencies are mitigated to some extent, thereby enhancing the reasoning performance. Evaluation on the benchmark method shows high linear relationship with model accuracy, implicating that additional module designed for reasoning order consistency may help boosting model performance.
It must be acknowledged that this work still possesses several limitations. Firstly, due to budget constraints, we were unable to conduct experiments on larger datasets or with more extensive large language models. Although we believe that the current experiments sufficiently demonstrate the viability and performance boost of the prompt strategy introduced in this research, additional experiments could potentially reveal more reliable patterns from the benchmark results. Furthermore, we lacked the computational resources and data necessary to retrain large language models. Originating from the hypotheses proposed in Section Methods, we posit that improvements could be made to large language models by enhancing the decoder part. Since the current output cannot see subsequent outputs, a potential research direction involves integrating the output layer with the original hidden state vectors before output layer, enabling the decoder to “see” the previous response during output. (Goyal et al. 2023) has already shown the potential to pause token to apply inference-time delay, introducing performance gain on various of reasoning datasets, showing that allowing for revisiting and revising its own answers may mitigate such issues. Therefore, a modified decoder network to explicitly combine possible future output may be a direct approach to mitigate hallucination in language models.